# CADDreamer: CAD object Generation from Single-view Images

### v1

@@ 4.1 Comparisons on Synthetic Images
Fig. 7 visually showcases the reconstruction results from synthetic images, while Tab. 2 and 3 provide quantitative comparisons. These results demonstrate that CADDreamer offers significant advantages in mesh generation quality, primitive extraction, and the topological fidelity of the reconstructed B-reps.
CADDreamer’s key strength lies in its ability to simultaneously capture low-level geometric features (normal maps) and high-level semantic shape understanding (primitive maps), resulting in meshes with minimal distortion. As shown in Fig. 7, our method generates clean, compact meshes with sharp edges, minimizing fitting errors and reconstruction failures. In contrast, competing methods, which lack high-level semantic understanding, generate distorted reconstructions that significantly deviate from the primitives. Quantitative results in Tab. 2 further validate CADDreamer’s performance, showing that its meshes exhibit the least geometric deviation from the ground truth, with the lowest CD and highest NC values.
In addition, CADDreamer employs a back-projection and Graph Cut process for primitive extraction, rather than relying on traditional mesh or point cloud segmentation methods. This approach enables more precise primitive extraction. Single-view RGB image reconstructions often produce meshes with errors such as uneven surfaces and over-smoothed boundaries. When these flawed meshes are processed through segmentation-based methods for primitive extraction, the reconstruction errors are amplified. For example, two twisted planes with a smooth transition might be misidentified as a single cone, as shown in Fig. 7(a, LRM). Using such incorrectly identified primitives for mesh reconstruction inevitably results in excessive fitting errors and reconstruction failures. In contrast, our method demonstrates superior primitive extraction that aligns closely with ground truth CAD models. Our segmentation approach achieves more precise primitive labeling than other methods, with $97.9\%$ of meshes correctly matching the primitive count of the ground truth, as shown in Tab. 2.
Thirdly, CADDreamer refines critical relationships between adjacent primitives, such as intersection, parallelism, perpendicularity, and collinearity. It also employs a topology-preserving intersection strategy to reconstruct B-reps with accurate topology, reducing two common errors: hanging faces that arise from incorrect primitive intersections, and massive non-closed B-reps caused by non-intersection, as shown in Fig. 7(a-f). As shown in Tab. 3, CADDreamer minimizes the percentages of hanging faces (HF) while significantly reducing face reconstruction failures caused by non-intersection, resulting in the lowest CD among compared approaches.
Figure 8: Segmentation results and reconstructed B-reps from real images (color definitions are the same as in Fig. 7).
@@ 4.2 Comparisons on Real Images
We captured RGB images of real-world CAD objects using a Canon EOS R5 camera with a Canon EF 70 lens. Era3D [20] was employed for background removal and as a normal predictor to generate accurate normal maps, which were subsequently used as inputs to evaluate our method’s generalization capability on real-world CAD objects. As shown in Fig. 8, despite the fact that real images possess more complex lighting and shadow effects, greater perspective distortion, and introduce larger errors in normal estimation, our method still successfully reconstructed high-quality CAD models, which demonstrates the generalization potential of our approach in handling real images.
@@ 5 Conclusion and Future Work
We intend to publicly release the source code and data once the paper is published. In this supplementary document, we provide more details about the evaluation metrics used in this paper, the ablation study, our limitation and failure cases, as well as more experimental results.
6 Details of Evaluation Metrics
@@
Table 4: Statistical results of reconstructed meshes and extracted primitives across ablation models. Evaluation metrics include: Chamfer distance (CD ×100), normal consistency (NC, %), vertex-based segmentation accuracy (SEG(V), %), and primitive count-based segmentation accuracy (SEG(P), %). Best values are highlighted in bold.
@@
Model 2 23.5 6.47
Model 3 11.8 3.29
@@
Table 5: Statistical results of reconstructed B-reps across ablation models, including the percentage of B-reps with hanging faces (HF, %), and Chamfer distance (CD ×$\times 100$) between the reconstructed B-reps and ground truth. Best values are highlighted in bold.
7 Ablation Study
@@
Our method benefits primarily from two distinct technical contributions: a cross-domain generation strategy that yields higher-quality meshes and improved segmentation results, and a geometric optimization strategy that reconstructs essential topological and geometric relationships to ensure CAD reconstruction through primitive intersection. We conduct ablation studies on these two technical contributions to investigate the performance gains achieved independently by each module.
First, we evaluate the contribution of our cross-domain generation strategy by replacing our multi-view generation module with original Wonder3D and Point2CAD, which we denote as Model 1. Tab. 4 presents the performance results of Model 1. While Wonder3D demonstrates improved generation results compared to previous methods with outputs closer to ground truth, non-negligible errors persist, with CD and NC metrics significantly trailing our approach. Moreover, these mesh reconstruction errors adversely affect Point2CAD’s segmentation performance and B-rep reconstruction accuracy, resulting in approximately 32.4% hanging faces, as detailed in Tab. 5.
Next, we investigate the contribution of our geometric optimization strategy and the importance of enforcing parallelism, perpendicularity, and collinearity constraints. We construct two variant models: (1) Model 2, which removes all geometric optimization steps from CADDreamer, and (2) Model 3, which maintains primitive stitching but eliminates the three geometric constraints. Since Models 2 and 3 share the same mesh reconstruction and segmentation pipeline with our method, we focus our analysis on comparing the reconstructed B-reps with ground truth, as shown in Tab. 5. Given that our method provides mesh and segmentation results closer to ground truth, both hanging faces and CD metrics of Models 2 and 3 demonstrate significant improvement over Model 1. However, without geometric optimization, reconstruction fitting errors introduce gaps between primitives and compromise their geometric relationships, leading to missing or incorrect intersections. Model 3, incorporating primitive stitching, effectively reduces intersection failures, thus achieving lower hanging face ratios and CD values compared to Model 2. Nevertheless, since Model 3 disregards geometric constraints, reconstruction failures caused by erroneous geometric relationships persist, resulting in more hanging faces and larger Chamfer distances than our CADDreamer.
@@
Figure 9: The limitations of our methods. (a) The multi-view images of six horizontal angles cannot cover the top face (red) of the CAD object. (b) An extremely thin and elongated plane (less than 1 pixel wide). (c) Our generated B-reps cannot preserve symmetry of the CAD structure.
8 Limitations and Failure Cases
@@
Constrained by limited multi-view angles and image resolutions, CADDreamer fails to maintain robustness under extreme viewing angles and struggles with CAD shapes containing fine geometric structures. First, CADDreamer generates mesh segmentation schemes using multi-view images, requiring the selected viewpoints to include all CAD faces. Since our method uses Wonder3D as a pretrained model with the same multi-view setup (six 256×256 images), our method requires that the CAD faces must be included in these six viewpoints. As illustrated in Fig. 9(a), when the six viewpoints used for generation are all horizontal, the top face of the CAD object does not appear in any of the six views. Consequently, the top face is omitted in the final generation, leading to inaccurate topology and reconstruction failure. Second, some B-reps contain extremely thin and elongated CAD faces that, when projected into multi-view images, manifest as sub-pixel features with widths less than one pixel. As demonstrated in Fig. 9(b), the elongated top face is segmented into isolated pixels by feature lines, which is not a continuous patch and cannot generate meaningful segmentation results. A potential solution to these limitations is to increase the number of views and improve image resolution: adding a top view ensures coverage of the top face, while using higher image resolution reduces the relative proportion of feature lines, enabling better face coverage. However, this approach demands prohibitive computational resources (requiring over 80GB of training GPU memory) and introduces greater convergence challenges.
@@
CADDreamer only reconstructs basic geometric relationships between primitives, such as parallelism, perpendicularity, and collinearity. It does not consider more complex relationships between multiple primitives, such as symmetry. As shown in Fig. 9(c), while the ground truth (upper) exhibits perfect symmetry between primitives, such property is difficult to be preserved in our results (lower) due to inevitable reconstruction and fitting errors. Furthermore, since detecting and optimizing such complex relationships remains an open problem, implementing their reconstruction within the geometric optimization module presents significant challenges.
9 More Experimental Results
@@
Fig. 1011, and 12 showcase more reconstruction results of our method, including input images, reconstructed meshes, reconstructed B-reps, as well as their CAD vertices and edges.
@@
Figure 10: Reconstruction results from the given images are shown from left to right: input image, reconstructed mesh, B-reps, and their CAD vertices and edges. Different colors indicate different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ).
@@
Figure 11: Reconstruction results from the given images are shown from left to right: input image, reconstructed mesh, B-reps, and their CAD vertices and edges. Different colors indicate different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ).
@@
Figure 12: Reconstruction results from the given images are shown from left to right: input image, reconstructed mesh, B-reps, and their CAD vertices and edges. Different colors indicate different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ).

### v2

@@ 4.1 Comparisons on Synthetic Images
Figure 6 visually showcases the reconstruction results from synthetic images, while Table 2 and 3 provide quantitative comparisons. These results demonstrate that CADDreamer offers significant advantages in mesh generation quality, primitive extraction, and the topological fidelity of the reconstructed B-reps.
CADDreamer’s key strength lies in its ability to simultaneously capture low-level geometric features (normal maps) and high-level semantic shape understanding (primitive maps), resulting in meshes with minimal distortion. As shown in Figure 6, our method generates clean, compact meshes with sharp edges, minimizing fitting errors and reconstruction failures. In contrast, competing methods, which lack high-level semantic understanding, generate distorted reconstructions that significantly deviate from the primitives. Quantitative results in Table 2 further validate CADDreamer’s performance, showing that its meshes exhibit the least geometric deviation from the ground truth, with the lowest CD and highest NC values.
In addition, CADDreamer employs a back-projection and Graph Cut process for primitive extraction, rather than relying on traditional mesh or point cloud segmentation methods. This approach enables more precise primitive extraction. Single-view RGB image reconstructions often produce meshes with errors such as uneven surfaces and over-smoothed boundaries. When these flawed meshes are processed through segmentation-based methods for primitive extraction, the reconstruction errors are amplified. For example, two twisted planes with a smooth transition might be misidentified as a single cone, as shown in Figure 6(a) LRM. Using such incorrectly identified primitives for mesh reconstruction inevitably results in excessive fitting errors and reconstruction failures. In contrast, our method demonstrates superior primitive extraction that aligns closely with ground truth CAD models. Our segmentation approach achieves more precise primitive labeling than other methods, with $97.9\%$ of meshes correctly matching the primitive count of the ground truth, as shown in Table 2.
Thirdly, CADDreamer refines critical relationships between adjacent primitives, such as intersection, parallelism, perpendicularity, and collinearity. It also employs a topology-preserving intersection strategy to reconstruct B-reps with accurate topology, reducing two common errors: hanging faces that arise from incorrect primitive intersections, and massive non-closed B-reps caused by non-intersection, as shown in Figure 6(a-f). As shown in Table 3, CADDreamer minimizes the percentages of hanging faces (HF) while significantly reducing face reconstruction failures caused by non-intersection, resulting in the lowest CD among compared approaches.
Figure 7: Segmentation results and reconstructed B-reps from real images (color definitions are the same as in Figure 6).
@@ 4.2 Comparisons on Real Images
We capture RGB images of real-world CAD objects using a Canon EOS R5 camera with a Canon EF 70 lens. Era3D [21] is employed for background removal and as a normal predictor to generate accurate normal maps, which are subsequently used as inputs to evaluate our method’s generalization capability on real-world CAD objects. As shown in Figure 7, despite the fact that real images possess more complex lighting and shadow effects, greater perspective distortion, and introduce larger errors in normal estimation, our method still successfully reconstructs high-quality CAD models, which demonstrates the generalization potential of our approach in handling real images.
@@ 5 Conclusion and Future Work
[…]
Figure A1: An illustration of the topology-preserving intersection strategy to achieve a watertight B-rep model with CAD vertices, edges, and faces. (a) depicts the topology derived from the mesh patches. (b) shows the CAD curves resulting from primitive intersections. (c) represents the reconstructed B-rep model through topology-preserving trimming.
Our project page and code can be found here. In this supplementary document, we provide more details about the evaluation metrics used in this paper, the ablation study, our limitations, and failure cases, as well as more experimental results.
Appendix A Details of Evaluation Metrics
@@
Table A1: Statistical results of reconstructed meshes and extracted primitives across ablation models. Evaluation metrics include: Chamfer distance (CD ×100), normal consistency (NC, %), vertex-based segmentation accuracy (SEG(V), %), and primitive count-based segmentation accuracy (SEG(P), %). Best values are highlighted in bold.
@@
Model 2 29.8 8.13
Model 3 2.40 1.37
Model 4 23.5 6.47
Model 5 11.8 3.29
@@
Table A2: Statistical results of reconstructed B-reps across ablation models, including the percentage of B-reps with hanging faces (HF, %), and Chamfer distance (CD ×$\times 100$) between the reconstructed B-reps and ground truth. Best values are highlighted in bold.
Appendix B Ablation Study
@@
Our method benefits primarily from two distinct technical contributions: a new multi-view generation module that yields higher-quality meshes and improved segmentation results, and a geometric optimization strategy that reconstructs essential topological and geometric relationships to ensure CAD reconstruction through primitive intersection. We conduct ablation studies on these two technical contributions to investigate the performance gains achieved independently by each module.
Figure A2: Using RGB images as input for fine-tuning will lead to poor generalization performance on real images. (a) real image input, (b-c) are the segmentation results of Model 2-3, which are fine-tuned with RGB inputs.
@@
Figure A3: The limitations of our methods. (a) The multi-view images of six horizontal angles cannot cover the top face (red) of the CAD object. (b) An extremely thin and elongated plane (less than 1 pixel wide). (c) A failure case of complex geometry (original image and segmentation result). (d) Our generated B-reps cannot preserve symmetry of the CAD structure.
Generation module. We first evaluate the contributions of our generation module through three comparative models. Model 1 employs the non-finetuned Wonder3D (RGB inputs and RGB&normal outputs) combined with Point2CAD to establish a baseline for comparison with other fine-tuned solutions. Model 2 utilizes the fine-tuned Wonder3D (RGB inputs and RGB&normal outputs) with Point2CAD to assess the effectiveness of our new cross-domain generation strategy (normal&semantic map). Model 3 implements CADDreamer’s pipeline with RGB inputs to evaluate the impact of our normal inputs. As demonstrated in Table A1, all models fine-tuned on primitive-based shapes outperform the original Wonder3D (Model 1), which lacks prior knowledge of primitives. Models that incorporate our cross-domain generation strategy for mesh segmentation (Models 3-5) achieve superior segmentation results compared to the post-processing segmentation approach applied to noisy meshes (Models 1-2), leading to more accurate B-reps reconstruction outcomes. Although using RGB images as inputs (Model 3) does not diminish CADDreamer’s performance on synthetic data, it impairs generalization to real-world image inputs, as illustrated in Figure A2. This limitation stems from the training data’s restricted texture variety and fixed lighting conditions, which fail to capture the complexity and variability of textures and illumination encountered in real-world scenarios.
Geometry optimization module. Next, we investigate the contribution of our geometric optimization strategy and the importance of enforcing parallelism, perpendicularity, and collinearity constraints. We construct two variant models: (1) Model 4, which removes all geometric optimization steps from CADDreamer, and (2) Model 5, which maintains primitive stitching but removes the three geometric constraints. Since Models 4 and 5 share the same mesh reconstruction and segmentation results with our method, we focus our analysis on comparing the reconstructed B-reps with ground truth, as shown in Table A2. Given that our method provides mesh and segmentation results closer to ground truth, the hanging faces and CD metrics of Models 4 and 5 demonstrate a significant improvement over Point2CAD (Model 1-2). Without geometric optimization, reconstruction fitting errors introduce gaps between primitives, compromising geometric relationships and causing missing or incorrect intersections (see Figure 4 in the main manuscript). Model 5, incorporating primitive stitching, effectively reduces intersection failures, thus achieving lower hanging face ratios and CD values compared to Model 4. Additionally, our model with three geometric constraints reduces reconstruction failures caused by erroneous geometric relationships, resulting in fewer hanging faces and smaller Chamfer distances than Model 5.
Appendix C Limitations and Failure Cases
CADDreamer consists of two main modules: the generation module and the geometry optimization module. We discuss the limitations of these two modules separately.
@@ Appendix C Limitations and Failure Cases
Generation module. Constrained by limited multi-view angles and image resolutions, the generation module fails to maintain robustness under extreme viewing angles and struggles with CAD shapes containing fine geometric structures. First, CADDreamer generates mesh segmentation schemes using multi-view images, requiring the selected viewpoints to include all CAD faces. Since our method uses Wonder3D as a pretrained model with the same multi-view setup (six 256×256 images), our method requires that the CAD faces must be included in these six viewpoints. As illustrated in Figure A3(a), when the six viewpoints used for generation are all horizontal, the top face of the CAD object does not appear in any of the six views. Consequently, the top face is omitted in the final generation, leading to inaccurate topology and reconstruction failure. Second, some B-reps contain extremely thin and elongated CAD faces that, when projected into multi-view images, manifest as sub-pixel features with widths less than one pixel. As demonstrated in Figure A3(b), the elongated top face is segmented into isolated pixels by feature lines, which is not a continuous patch and cannot generate meaningful segmentation results. The combination of these limitations means that CADDreamer might fail to reconstruct and segment complex CAD shapes, especially shapes with fine structures and excessive occlusion, as shown in Figure A3(c). A potential solution to these limitations is to increase the number of views and improve image resolution: adding a top view ensures coverage of the top face, while using higher image resolution makes fine structures more visible, enabling better face coverage. However, this approach demands prohibitive computational resources and introduces greater convergence challenges.
@@
Geometry optimization module only reconstructs basic geometric relationships between primitives, such as parallelism, perpendicularity, and collinearity. It does not consider more complex relationships between multiple primitives, such as symmetry. As shown in Figure A3(d), while the ground truth (upper) exhibits perfect symmetry between primitives, such property is difficult to be preserved in our results (lower) due to inevitable reconstruction and fitting errors. Furthermore, since detecting and optimizing such complex relationships remains an open problem, implementing their reconstruction within the geometric optimization module presents significant challenges.
Appendix D More Comparison Results
We add additional comparative experiments to analyze the advantages of CADDreamer. First, we add the comparison with DeepCAD [41] to analyze the advantages of CADDreamer compared to sketch-extrude methods. In addition, we notice some closed-source (Tripo [38]) and open-source (Trellis [42], MeshAnything [5]) models released concurrently with CADDreamer. We analyze the performance of these methods on some typical examples and explain why these methods are not suitable for CAD generation.
Figure A4: Comparison with other methods. From left to right, (a) Input image, (b) DeepCAD results [41], (c) Tripo results [38], (d) Trellis results [42], (e) MeshAnything results [5]), and (f) our segmentation results. Color definitions match those used in the main manuscript figures.
Since DeepCAD cannot handle the image inputs, we use our generated meshes as inputs and use DeepCAD to generate B-reps. As shown in Figure A4, since sketch-extrude approaches can only support planes and cylinders, complex primitives like cones are simplified, and the reconstructed B-reps have higher errors (CD=4.21, NC=67.8) than CADDreamer (CD=1.36, NC=91.9) on our testing dataset.
We also evaluate Tripo, Trellis, and MeshAnything on 10 diverse CAD models for mesh reconstruction, followed by Point2CAD for CAD generation to obtain CAD models. Tripo and Trellis tend to add unnecessary details and components due to their focus on generating diverse and complex shapes, which results in a large number of segmented primitives and faces, as shown in Figure A4. MeshAnything, based on a finite-length autoregressive generation progress, overlooks strong constraints on mesh connectivity. Consequently, it tends to generate simple and topology-breaking meshes, resulting in numerous isolated primitives and inferior outcomes than CADDreamer.
Appendix E More Reconstruction Results
@@
Figure A5A6, and A7 showcase more reconstruction results of our method, including input images, reconstructed meshes, reconstructed B-reps, as well as their CAD vertices and edges.
@@
Figure A5: Reconstruction results from the given images are shown from left to right: input image, reconstructed mesh, B-reps, and their CAD vertices and edges. Different colors indicate different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ).
@@
Figure A6: Reconstruction results from the given images are shown from left to right: input image, reconstructed mesh, B-reps, and their CAD vertices and edges. Different colors indicate different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ).
@@
Figure A7: Reconstruction results from the given images are shown from left to right: input image, reconstructed mesh, B-reps, and their CAD vertices and edges. Different colors indicate different primitives: plane ( ), cylinder ( ), cone ( ), sphere ( ), and torus ( ).